# A Review On Safe Reinforcement Learning Using Lyapunov and Barrier Functions

### v2

@@ 4.2 Barrier Functions for Reinforcement Learning
Lyapunov and Barrier Methods Lyapunov-Barrier Function [109, 110, 111, 112, 113, 114, 115] Lyapunov-Barrier-based Reward [116, 115] Lyapunov-Barrier-based Loss [124, 123, 111] Lyapunov-Barrier-RL-QP [117, 118, 119, 120, 122, 122]
@@
4.3 Lyapunov and Barrier Functions for Reinforcement Learning
Few recent approaches have leveraged the stability guarantees provided by Lyapunov functions and the constraint satisfaction capability of barrier functions to form a single approach. Fig. 6 shows a broad classification of approaches using Lyapunov and barrier methods. In [117], the authors propose use of temporal logic based RL using CBFs for the PPO algorithm in a dynamic obstacle avoidance problem. The authors consider a syntactically co-safe truncated temporal logic which is used to generate a finite state automata (FSA), this generated FSA is then augmented with the MDP to introduce temporal logic in the RL formulation. A CLF-CBF-QP (Proposition 1 in Section III of cited work) essentially formulates the CLF and CBF as constraints for a QP problem with a quadratic cost as the safe action. This formulated QP is relaxed using $\delta$ to avoid deadlock between CLF and CBF constraints. The CLF-CBF-QP for a general continuous time nonlinear control affine dynamical system is given as follows:
@@ 4.3 Lyapunov and Barrier Functions for Reinforcement Learning
Li et al.(2019) [117] Temporal logic guided CLF-CBF-QP based policy optimization PPO Unicycle
Choi et al. (2020) [118] CLF-CBF-QP based policy optimization DDPG Bipedal robot
@@
Du et al. (2023) [109] Control Lyapunov-Barrier function as value function Actor-critic 2D quadrotor
Cohen et al. (2023) [110] Lyapunov-like CBF ADP Collision Avoidance
Zhao et al. (2023) [121] CLF and CBF constraints as a Lagrangian loss function Actor-critic Unicycle
Meng et al. (2023) [119] CLF-control dependent CBF-time varying CBF-RL-QP optimization. CLF, CDBF and TCBF are the constraints for RL QP formulations DDPG CarSim
Jang et al. (2024) [120] Lyapunov-based CBF in a QP formulation LQR Multirotor, Cartpole, ground rover
Wei et al. (2024) [113] Asymmetric time-varying integral barrier Lyapunov function in optimized backstepping control Actor-critic Pendulum, one-link robot
Mizuta et al. (2024) [116] CLF and CBF based reward functions Diffusion model Custom
Zhao et al. (2024) [111] CLF and CBF constraints as a Lagrangian loss function SAC Simulated car following
Wang et al. (2024) [115] Control Lyapunov Barrier function augmented rewards Policy Iteration Chemical Process
Mandal et al. (2024) [112] Formal verification on neural Lyapunov barrier certificates PPO Custom - Spacecraft control
Chen et al. (2024) [122] CLF-CBF-QP based low level controller. DQN based high level controller. DQN Simulink - autonomous driving
Cocaul et al. (2024) [123] Empirical Lyapunov risk term and barrier in critic loss function. PPO OpenAI Gym
Zhao et al. (2025) [124] Lagrangian-based Lyapunov and barrier constraints augmented with policy loss function. Actor-critic OpenAI Gym, Safe-Control-Gym
@@
The approach combines model-based planning (temporal logic) and model-free control (RL). Choi et al. [118] use RL to learn model uncertainty for CLF and CBF dynamic constraints for nonlinear bipedal robot (assumption: input-output linearization) to address the issue of model uncertainty in data-driven RL. A QP formulation with CLF and CBF constraints (similar to [117]) is used to solve for the control. DDPG agent is used for learning uncertainty terms in CLF, CBF and other dynamic constraints (policy for RL). Authors in [119] follow a similar CLF-CBF-RL-QP formulation by considering CLF, control dependent (CDBF) and time-varying CBF (TCBF). In [120], propose Lyapunov-based CBF constraints for a QP formulation. Thus, variations of CLF-CBF-RL-QP are observed in literature to guarantee stability and satisfy state constraints. Mandal et al. [112] propose a framework that verifies neural Lyapunov barrier (NLB) certificates using a counterexample-guided interactive synthesis (CEGIS) loop [125] to obtain the RL controller. CEGIS loop augments the training dataset in case the policy does not satisfy the constraints of the certificate. Reach-while-avoid certificates are used as the NLBs with further extensions proposed in terms of filtering and compositions. The extensions essentially filters the state space according to safe and unsafe regions, and compositions help in scalability of filtering based approach by parallel training of multiple controllers which satisfy safe operation and selecting appropriate safe controller at a given state.
@@
Compared to prior approaches where a CLF-CBF-QP is solved to compensate for RL control output, Zhang et al. [114] propose a Barrier Lyapunov function-based RL approach for tracking and planning of vehicle motion control using an ADP algorithm. The unknown optimal policy is reformulated using Barrier-Lyapunov functions method into model-based and adaptive parts. The actor-critic networks are used to approximate the control policy and value function. Backstepping is employed to have global control of the system by optimizing virtual control in each subsystem. As in the case of prior approaches, backstepping is introduced to approximate the time derivative of along the dynamics for CLF and CBF. Du et al. [109], propose a purely data-driven approach using a control Lyapunov Barrier function in an actor-critic formulation for 2D quadrotor navigation. The control Lyapunov Barrier function (CLBF) is chosen as the critic function in a CMDP formulation with constraints based on data-based CLBF theorem (Section IV-B of cited work). The constrained optimization problem is augmented to an unconstrained problem using Lagrange multipliers and updated through stochastic gradient descent as follows,
@@
$-Q_{\text{LB}}(x,\pi_{\theta}(x))\mathbbm{1}_{\Delta}(x)+\alpha d\bigg{)}\bigg{]}$ (24)
where, $Q_{\text{LB}}$ is the Lyapunov-barrier function, $\mathbbm{1}_{\Delta}(\cdot)$ denotes the indicator function with respect to state $x$ being in the safe set ($\Delta=\mathcal{X}\backslash\mathcal{X}_{\text{goal}}\cup\mathcal{X}_{\text{unsafe}}$), $\lambda$ denotes the Lagrange multiplier, $\alpha$ is a positive constant and $d$ denotes the upper bound on constraint cost in a CMDP. Thus, the proposed $Q_{\text{LB}}$ satisfies the system and state constraints by satisfying the CMDP cost constraint at each step. Cohen et al. [110] follow a similar approach of using a Lyapunov-like CBF for a collision avoidance application in a model-based ADP formulation. The Lyapunov-like CBF is expressed as a re-centered barrier function [126],
@@
which gives suitable Lyapunov function properties of positive semi-definiteness. The barrier function is re-centered at the origin, thus, the barrier function vanishes at the origin. The proposed Lyapunov CBFs are added as a compensating controller (similar to 20) in the control policy to drive the control input to safe sets and guarantee stability. The problem is formulated as an ADP and tested on collision avoidance and single integrator system with unknown drift dynamics. Zhao et al. [121] propose a barrier-Lyapunov actor-critic approach using a Lagrangian formulation for the actor and critic loss functions. A backup controller is proposed in case valid control input can not be computed by the actor-critic RL controller ensuring safety and stability constraints. The authors also propose in [111] a neural ordinary differential equation (NODE)-based framework where NODE is used to predict the system dynamics. The remaining framework remains the similar to their prior work using the barrier-Lyapunov actor-critic approach. Authors in [123, 124] follow a similar methodology with [123] incorporating the empirical Lyapunov risk term in the critic loss function.
@@
Few approaches consider Lyapunov-barrier-based reward-shaping. Notably, Mizuta et al. [116] propose CLF and CBF terms to guide the denoising process of a diffusion model. CLF and CBF rewards functions are proposed for a single and multi-agent pedestrian dataset. Since we focus on safe RL in this review, a detailed explanation is omitted and the idea to introduce this work is to provide possible research directions for using reward-shaping.
Table 3 summarizes the literature reviewed for cumulative use of Lyapunov and barrier functions in RL. Most methods combining Lyapunov and barrier functions employ two of the following methods:(i) Solving a constrained optimization problem with Lyapunov and barrier constraints; (ii) using Lyapunov-barrier terms in the critic/actor loss function; (iii) Lyapunov-barrier-based reward shaping. It remains an open question for the feasibility and complexity of solving a constrained optimization problem at each iteration whilst optimizing the parameters for RL in case of large-scale problems, to satisfy the cumulative Lyapunov-Barrier constraints. A similar trend can also be observed in terms of selection for a candidate Lyapunov function for model-free approaches. Proving the effectiveness of approximating a candidate Lyapunov function with sampled data from the environment remains an open area of research.

### v3

@@ 5 Barrier Functions for Reinforcement Learning
5.1 Classification and Discussion
Table 4 provides a tabular comparison of each methodology and its shortcomings. Barrier-function-based safe RL methods have grown rapidly due to their ability to formalize safety as forward invariance of a set, i.e., trajectories remain in a constraint-admissible region when a control barrier function (CBF) condition is enforced. The surveyed literature can be organized around how the barrier enters the learning loop: (i) as a soft learning signal (reward/cost shaping), (ii) as a hard execution-time filter (CBF-QP shielding), and (iii) as a constraint embedded into policy optimization (Lagrangian/learner–verifier/bilevel formulations). These categories provide complementary trade-offs between scalability, theoretical rigor, and practical deployability.
Barrier shaping: simplicity with limited guarantee strength
Reward shaping with barrier terms modifies the cost-to-go or reward to discourage constraint violation while preserving performance objectives [marvi2021safe, liu2023safe2, ranjan2024barrier, liu2025safe, xie2025certificated]. This approach is attractive because it integrates seamlessly with standard actor–critic or policy iteration pipelines and can improve learning efficiency by biasing exploration away from unsafe states. However, shaping is typically a soft mechanism: unless the shaped objective is paired with additional constraint-enforcement machinery, safety violations can still occur particularly during early exploration or under distribution shift. Moreover, barrier weights introduce a brittle safety–optimality trade-off; poorly tuned penalties can either fail to prevent violations or over-regularize and yield conservative policies.
CBF shielding via QP filters: hard constraints with model dependence and feasibility challenges
A dominant template in barrier-based safe RL is to treat the RL policy as a nominal controller $u^{RL}$ and compute a minimally modified safe action via a CBF-QP filter [cheng2019end, emam2021safe, huang2022barrier, song2022safe, cheng2023safe, yang2024enhancing, dinh2024towards, kalaria2024disturbance, hou2024safe]. This “safety layer” can provide strong practical safety because it enforces barrier constraints at execution time whenever the QP is feasible. The main limitation is that computing the CBF constraints requires system information typically a control-affine structure and Lie derivative estimates, so most methods introduce a model learner (e.g., GP) or an estimator/observer to approximate unknown dynamics [emam2021safe, cheng2023safe, kalaria2024disturbance]. This creates a tight coupling between certificate validity and modeling accuracy. In addition, QP feasibility (and deadlock near boundaries) remains a core issue. Slack-variable relaxations address feasibility [wang2022ensuring], but transform strict safety into a penalty violation term that must be tuned and monitored.
Constrained policy optimization and verification: end-to-end training vs enforceability
Another class of approaches incorporates barrier constraints directly into the policy update, often through Lagrangian formulations or learner–verifier pipelines [zhao2022safe, zhang2022barrier, liu2023safe1, yang2023model, dey2024p2bpo, zhang2024constrained]. These methods can reduce reliance on per-step QPs and enable differentiable end-to-end training, but they frequently enforce safety in expectation or via approximate penalty mechanisms. Formal verification steps can strengthen guarantees [zhao2022safe], yet they introduce computational overhead and typically depend on the tractability of verifying barrier certificates. In stochastic environments, bilevel and chance-constrained formulations seek to capture probabilistic safety more directly [wang2023enforcing, sabouni2024reinforcement], but at the cost of substantial computational complexity and sensitivity to probabilistic model calibration.
Multi-agent and disturbance-aware settings: realistic but significantly harder
Realistic deployment often entails disturbances, unmodeled dynamics, and multi-agent coupling. Disturbance-aware CBF filters that leverage GPs or disturbance observers aim to make invariance robust to perturbations [emam2021safe, cheng2023safe, hou2024safe, kalaria2024disturbance]. Meanwhile, decentralized CBF shielding extends the QP-filter template to multi-agent collision avoidance [cai2021safe]. These extensions improve realism but amplify fundamental difficulties: estimating Lie derivatives reliably under uncertainty, maintaining feasibility under coupled constraints, and ensuring stability of the combined estimator–filter–RL closed loop.
5.2 Key Issues and Open Problems
Model dependence and certificate validity under approximation
Most CBF guarantees hinge on accurate Lie derivative computation. When dynamics are unknown, learned models/observers introduce approximation error that can invalidate safety certificates. A core research need is conservative-yet-nontrivial uncertainty bounds for learned derivatives that scale to high dimensions, enabling certificates that remain valid under model error.
Feasibility, deadlock, and principled slack allocation
QP filters can become infeasible near constraint boundaries or under conflicting constraints, causing deadlock. Slack variables improve feasibility [wang2022ensuring], but selecting slack penalties is ad hoc and may conceal systematic model error. Developing principled violation penalty terms (state-dependent slack penalties, risk-aware slack allocation, and safety margin adaptation) is a key direction.
Exploration under hard safety constraints
CBF shielding prevents unsafe actions but can severely restrict exploration, especially when the safe set is small or conservative due to modeling bounds. Methods that actively expand the safe set while exploring (e.g., safe set enlargement strategies, optimistic exploration within certified regions, or barrier-aware intrinsic motivation) remain underdeveloped compared to Lyapunov ROA expansion.
End-to-end differentiable safety layers and learning stability
Integrating QP filters into backpropagation (differentiable QPs) can reduce projection-induced bias and improve sample efficiency, but raises stability and computational challenges. Understanding how safety layers affect policy gradients, convergence, and optimality gaps is an open problem.
Chance constraints, distribution shift, and calibration
Probabilistic safety is essential in stochastic environments, but chance-constrained/bilevel approaches depend on calibrated uncertainty models [wang2023enforcing]. Research is needed on robust chance constraints under miscalibration, offline-to-online transfer, and rare-event safety validation.
Multi-agent coupled constraints and partial observability
Decentralized CBF shields [cai2021safe] are promising, yet coupled constraints yield complex feasibility regions and non-stationary learning dynamics. Open directions include compositional multi-agent barriers, scalable coordination protocols, and barrier certificates under partial observability and intermittent communication.
Barrier-function-based safe RL offers a pragmatic path to enforcing state constraints through forward invariance, with CBF-QP shielding emerging as a dominant tool. Table 3 summarizes the literature reviewed for cumulative use of Lyapunov and barrier functions in RL. The central trade-off mirrors Lyapunov-based methods: strong safety enforcement often requires substantial system knowledge and can induce conservatism, while purely learning-based barrier objectives scale more easily but provide weaker guarantees. Bridging this gap via scalable uncertainty quantification, feasibility-aware filters, and exploration that expands certified safe sets appears to be the key to widely deployable barrier-based safe RL.
Lyapunov and Barrier Methods Lyapunov-Barrier Function [du2023reinforcement, cohen2023safe, zhao2024nlbac, mandal2024formally, wei2024asymmetric, 9827972, wang2024control] Lyapunov-Barrier-based Reward [mizuta2024cobl, wang2024control] Lyapunov-Barrier-based Loss [zhao2025nlbac, cocaul2024safe, zhao2024nlbac] Lyapunov-Barrier-RL-QP [li2019temporal, choi2020reinforcement, meng2023reinforcement, jang2024safe, chen2024collision, chen2024collision]
@@ 5.2 Key Issues and Open Problems
6 Lyapunov and Barrier Functions for Reinforcement Learning
Few recent approaches have leveraged the stability guarantees provided by Lyapunov functions and the constraint satisfaction capability of barrier functions to form a single approach. Fig. 6 shows a broad classification of approaches using Lyapunov and barrier methods. In [li2019temporal], the authors propose use of temporal logic based RL using CBFs for the PPO algorithm in a dynamic obstacle avoidance problem. The authors consider a syntactically co-safe truncated temporal logic which is used to generate a finite state automata (FSA), this generated FSA is then augmented with the MDP to introduce temporal logic in the RL formulation. A CLF-CBF-QP (Proposition 1 in Section III of cited work) essentially formulates the CLF and CBF as constraints for a QP problem with a quadratic cost as the safe action. This formulated QP is relaxed using $\delta$ to avoid deadlock between CLF and CBF constraints. The CLF-CBF-QP for a general continuous time nonlinear control affine dynamical system is given as follows:
@@ 6 Lyapunov and Barrier Functions for Reinforcement Learning
Li et al.(2019) [li2019temporal] Temporal logic guided CLF-CBF-QP based policy optimization PPO Unicycle
Choi et al. (2020) [choi2020reinforcement] CLF-CBF-QP based policy optimization DDPG Bipedal robot
@@
Du et al. (2023) [du2023reinforcement] Control Lyapunov-Barrier function as value function Actor-critic 2D quadrotor
Cohen et al. (2023) [cohen2023safe] Lyapunov-like CBF ADP Collision Avoidance
Zhao et al. (2023) [zhao2023stable] CLF and CBF constraints as a Lagrangian loss function Actor-critic Unicycle
Meng et al. (2023) [meng2023reinforcement] CLF-control dependent CBF-time varying CBF-RL-QP optimization. CLF, CDBF and TCBF are the constraints for RL QP formulations DDPG CarSim
Jang et al. (2024) [jang2024safe] Lyapunov-based CBF in a QP formulation LQR Multirotor, Cartpole, ground rover
Wei et al. (2024) [wei2024asymmetric] Asymmetric time-varying integral barrier Lyapunov function in optimized backstepping control Actor-critic Pendulum, one-link robot
Mizuta et al. (2024) [mizuta2024cobl] CLF and CBF based reward functions Diffusion model Custom
Zhao et al. (2024) [zhao2024nlbac] CLF and CBF constraints as a Lagrangian loss function SAC Simulated car following
Wang et al. (2024) [wang2024control] Control Lyapunov Barrier function augmented rewards Policy Iteration Chemical Process
Mandal et al. (2024) [mandal2024formally] Formal verification on neural Lyapunov barrier certificates PPO Custom - Spacecraft control
Chen et al. (2024) [chen2024collision] CLF-CBF-QP based low level controller. DQN based high level controller. DQN Simulink - autonomous driving
Cocaul et al. (2024) [cocaul2024safe] Empirical Lyapunov risk term and barrier in critic loss function. PPO OpenAI Gym
Zhao et al. (2025) [zhao2025nlbac] Lagrangian-based Lyapunov and barrier constraints augmented with policy loss function. Actor-critic OpenAI Gym, Safe-Control-Gym
@@
The approach combines model-based planning (temporal logic) and model-free control (RL). Choi et al. [choi2020reinforcement] use RL to learn model uncertainty for CLF and CBF dynamic constraints for nonlinear bipedal robot (assumption: input-output linearization) to address the issue of model uncertainty in data-driven RL. A QP formulation with CLF and CBF constraints (similar to [li2019temporal]) is used to solve for the control. DDPG agent is used for learning uncertainty terms in CLF, CBF and other dynamic constraints (policy for RL). Authors in [meng2023reinforcement] follow a similar CLF-CBF-RL-QP formulation by considering CLF, control dependent (CDBF) and time-varying CBF (TCBF). In [jang2024safe], propose Lyapunov-based CBF constraints for a QP formulation. Thus, variations of CLF-CBF-RL-QP are observed in literature to guarantee stability and satisfy state constraints. Mandal et al. [mandal2024formally] propose a framework that verifies neural Lyapunov barrier (NLB) certificates using a counterexample-guided interactive synthesis (CEGIS) loop [katz2019marabou] to obtain the RL controller. CEGIS loop augments the training dataset in case the policy does not satisfy the constraints of the certificate. Reach-while-avoid certificates are used as the NLBs with further extensions proposed in terms of filtering and compositions. The extensions essentially filters the state space according to safe and unsafe regions, and compositions help in scalability of filtering based approach by parallel training of multiple controllers which satisfy safe operation and selecting appropriate safe controller at a given state.
Table 6: Classification combined Lyapunov–barrier (CLF/CBF/CLBF) approaches in safe RL.
Type Methodology Pros Cons
CLF–CBF–QP shielding / filtering Nominal $u_{RL}$ is corrected by solving a QP with CLF and CBF constraints (often relaxed by $\delta$); executed action is $u=u_{RL}+u^{\text{CLF-CBF}}$ [li2019temporal, choi2020reinforcement, meng2023reinforcement, jang2024safe]. Enforces stability + safety in one layer; plug-and-play with PPO/DDPG/SAC; prevents catastrophic rollouts when feasible; can constrain exploration to safe sets. CLF/CBF conflicts near boundaries $\Rightarrow$ slack weakens guarantees; per-step QP cost and poor scaling with many constraints; needs control-affine model/Lie derivatives (model/estimator dependence); action filtering biases gradients and off-policy credit.
Temporal-logic + CLF–CBF filtering Task specification via FSA/temporal logic augments MDP rewards; CLF–CBF-QP provides low-level safety and stabilization [li2019temporal]. Supports sequencing/reach–avoid specs; clean hierarchy: high-level task, low-level safety/stability; retains continuous-time constraints. Automata/product-MDP overhead; higher sample/compute cost; temporal goals may conflict with continuous safety; limited scalability for long-horizon specs.
Uncertainty-aware CLF/CBF constraints RL learns model uncertainty terms that enter CLF/CBF constraints; QP enforces constraints with learned uncertainty compensation [choi2020reinforcement]. Handles model mismatch; less conservative than worst-case bounds; improves robustness in nonlinear robotics. Uncertainty miscalibration under shift can break safety; coupled learning+constraints can destabilize; extra tuning and components.
Enhanced barrier classes within CLF–CBF-QP (CDBF/TCBF, Lyapunov-based CBF) Extend CBF constraints (control-dependent, time-varying, Lyapunov-based CBF forms) while retaining CLF constraints in RL-QP [meng2023reinforcement, jang2024safe]. Better for time-varying/dynamic constraints; can improve feasibility and reduce conservatism; richer safety modeling. More parameters/constraints $\Rightarrow$ harder tuning + worse scaling; needs accurate timing/state; certification assumptions become heavier.
Verification-driven neural Lyapunov barrier (NLB) certificates Train policy and neural certificates with formal verification (CEGIS); counterexamples augment data until certificate constraints hold [mandal2024formally]. Stronger assurance than empirical training; counterexamples expose failure modes; supports reach-while-avoid certificates. High verification cost; hard to scale CEGIS to high-dimensional control; guarantees depend on verifier + certificate class; many iterations may be needed.
Barrier-Lyapunov in ADP / backstepping Barrier-Lyapunov reformulates optimal policy into model-based + adaptive parts; actor–critic approximates policy/value; backstepping handles nonlinearities/derivatives [9827972, wei2024asymmetric]. Structured nonlinear design; systematic subsystem/virtual-control handling; interpretable stability/constraint mechanism. Model/derivative dependence; high engineering complexity; limited portability beyond assumed structure; can be conservative.
CLBF / Lyapunov–barrier critic in CMDP actor–critic Single Lyapunov–barrier object (e.g., $Q_{\mathrm{LB}}$) used as critic; Lagrangian enforces CMDP constraint costs on safe set indicator [du2023reinforcement]. No per-step QP; end-to-end learning with safety in critic; CMDP-compatible and data-driven; simple deployment. CLBF conditions are hard to satisfy under function approximation/shift; multiplier tuning is brittle; constraints often approximate unless paired with filters.
Lyapunov-like CBF via re-centered barrier (ADP) Construct Lyapunov-like barrier $H(x)=(h(x)-h(0))^{2}$ to obtain positive semi-definite Lyapunov properties; used in ADP with compensating control [cohen2023safe, wills2004barrier]. Unifies Lyapunov positivity with barrier geometry; interpretable (vanishes at origin); can stabilize while avoiding unsafe sets. Sensitive to choice of $h(\cdot)$ and model/derivatives; limited for complex safe sets; feasibility/optimization issues persist in filtering.
Barrier-Lyapunov actor–critic with backup policies / NODE dynamics Actor–critic losses augmented with Lyapunov+barrier constraints; backup controller used if learned policy unsafe; NODE predicts dynamics to support constraints [zhao2023stable, zhao2024nlbac, zhao2025nlbac, cocaul2024safe]. Practical safety via fallback; NODE reduces mismatch for constraint evaluation; flexible in deep RL; risk terms can shape learning. Safety hinges on learned dynamics/certificates + backup quality; verification remains difficult; more components $\Rightarrow$ more failure modes; weight/multiplier tuning is sensitive.
Lyapunov–barrier reward shaping (emerging) CLF/CBF terms added to rewards to guide learning (including diffusion-model guided control) [mizuta2024cobl]. Easy to add; encourages safe/stable behavior; can improve exploration and sample efficiency. Soft safety only (no invariance); shaping can distort optimality; limited hard-safety evidence in benchmarks; theory for joint shaping remains sparse.
@@
Compared to prior approaches where a CLF-CBF-QP is solved to compensate for RL control output, Zhang et al. [9827972] propose a Barrier Lyapunov function-based RL approach for tracking and planning of vehicle motion control using an ADP algorithm. The unknown optimal policy is reformulated using Barrier-Lyapunov functions method into model-based and adaptive parts. The actor-critic networks are used to approximate the control policy and value function. Backstepping is employed to have global control of the system by optimizing virtual control in each subsystem. As in the case of prior approaches, backstepping is introduced to approximate the time derivative of along the dynamics for CLF and CBF. Du et al. [du2023reinforcement], propose a purely data-driven approach using a control Lyapunov Barrier function in an actor-critic formulation for 2D quadrotor navigation. The control Lyapunov Barrier function (CLBF) is chosen as the critic function in a CMDP formulation with constraints based on data-based CLBF theorem (Section IV-B of cited work). The constrained optimization problem is augmented to an unconstrained problem using Lagrange multipliers and updated through stochastic gradient descent as follows,
@@
$-Q_{\text{LB}}(x,\pi_{\theta}(x))\mathbbm{1}_{\Delta}(x)+\alpha d\bigg)\bigg]$
where, $Q_{\text{LB}}$ is the Lyapunov-barrier function, $\mathbbm{1}_{\Delta}(\cdot)$ denotes the indicator function with respect to state $x$ being in the safe set ($\Delta=\mathcal{X}\backslash\mathcal{X}_{\text{goal}}\cup\mathcal{X}_{\text{unsafe}}$), $\lambda$ denotes the Lagrange multiplier, $\alpha$ is a positive constant and $d$ denotes the upper bound on constraint cost in a CMDP. Thus, the proposed $Q_{\text{LB}}$ satisfies the system and state constraints by satisfying the CMDP cost constraint at each step. Cohen et al. [cohen2023safe] follow a similar approach of using a Lyapunov-like CBF for a collision avoidance application in a model-based ADP formulation. The Lyapunov-like CBF is expressed as a re-centered barrier function [wills2004barrier],
@@
which gives suitable Lyapunov function properties of positive semi-definiteness. The barrier function is re-centered at the origin, thus, the barrier function vanishes at the origin. The proposed Lyapunov CBFs are added as a compensating controller (similar to 20) in the control policy to drive the control input to safe sets and guarantee stability. The problem is formulated as an ADP and tested on collision avoidance and single integrator system with unknown drift dynamics. Zhao et al. [zhao2023stable] propose a barrier-Lyapunov actor-critic approach using a Lagrangian formulation for the actor and critic loss functions. A backup controller is proposed in case valid control input can not be computed by the actor-critic RL controller ensuring safety and stability constraints. The authors also propose in [zhao2024nlbac] a neural ordinary differential equation (NODE)-based framework where NODE is used to predict the system dynamics. The remaining framework remains the similar to their prior work using the barrier-Lyapunov actor-critic approach. Authors in [cocaul2024safe, zhao2025nlbac] follow a similar methodology with [cocaul2024safe] incorporating the empirical Lyapunov risk term in the critic loss function.
@@
Few approaches consider Lyapunov-barrier-based reward-shaping. Notably, Mizuta et al. [mizuta2024cobl] propose CLF and CBF terms to guide the denoising process of a diffusion model. CLF and CBF rewards functions are proposed for a single and multi-agent pedestrian dataset. Since we focus on safe RL in this review, a detailed explanation is omitted and the idea to introduce this work is to provide possible research directions for using reward-shaping.
6.1 Classification and Discussion
Table 6 provides a tabular comparison of each methodology and its shortcomings. The surveyed literature can be broadly grouped into three recurring templates: (i) CLF-CBF constrained optimization filters, where an RL action is compensated via a per-step quadratic program (QP) constrained by both Lyapunov and barrier conditions; (ii) Lyapunov-barrier critics and Lagrangian learning, where a single Lyapunov-barrier function (e.g., CLBF, Lyapunov-like CBF) is embedded into the critic or actor-critic objective; and (iii) Lyapunov–barrier reward shaping, where CLF/CBF terms shape rewards to bias exploration toward stability and constraint satisfaction.
CLF-CBF constrained optimization filters
A dominant approach is to treat the RL controller as a nominal policy $u_{RL}$ and compute the final action using a CLF–CBF-QP safety filter. In [li2019temporal], temporal-logic-guided RL augments the MDP with a finite-state automaton and solves a relaxed CLF–CBF-QP to prevent deadlock between stabilization and constraint satisfaction. The resulting control law is of the form $u_{k}(x)=u_{k}^{RL}+u_{k}^{\text{CLF-CBF}}(x,u_{k}^{RL})$, where the compensating term is obtained by enforcing Lyapunov decrease and barrier constraints. Variants of this paradigm also appear in [choi2020reinforcement, meng2023reinforcement, jang2024safe], differing primarily in (i) how uncertainty/disturbances are learned or estimated (e.g., learning uncertainty terms in constraints [choi2020reinforcement]), and (ii) the choice of barrier type (e.g., control-dependent and time-varying CBFs [meng2023reinforcement]) or Lyapunov-based barrier formulations [jang2024safe]. While these filters offer strong practical safety when feasible, they raise important questions regarding feasibility under simultaneous CLF and CBF constraints, the computational cost of solving a constrained optimization at each time step, and the induced conservatism that can limit exploration.
Lyapunov-barrier critics and Lagrangian formulations
A second class of methods integrates Lyapunov–barrier structure directly into the learning objective to avoid (or complement) per-step QPs. Zhang et al. [9827972] employ barrier-Lyapunov formulations in an ADP setting for vehicle tracking, using actor–critic approximations and backstepping to handle nonlinearities and derivative terms. Du et al. [du2023reinforcement] propose a data-driven control Lyapunov–barrier function (CLBF) as the critic in a CMDP-based actor–critic formulation, using Lagrangian multipliers to enforce constraint costs on the learned Lyapunov–barrier critic. Cohen et al. [cohen2023safe] similarly leverage a Lyapunov-like CBF via a re-centered barrier construction to obtain positive semi-definite Lyapunov properties, embedding the resulting certificate in an ADP framework for collision avoidance. Zhao et al. [zhao2023stable] propose a barrier-Lyapunov actor–critic method with Lagrangian losses and incorporate a backup controller when the learned policy cannot provide valid safe actions; follow-up work introduces NODE-based dynamics prediction to support the same learning framework [zhao2024nlbac], with related extensions including empirical Lyapunov risk in critic learning [cocaul2024safe, zhao2025nlbac]. These methods reduce reliance on online QP solving but shift the main challenge to reliably learning certificates from sampled data and ensuring generalization beyond the training distribution.
Formal verification and certificate synthesis
Mandal et al. [mandal2024formally] propose verifying neural Lyapunov barrier (NLB) certificates using a counterexample-guided interactive synthesis (CEGIS) loop [katz2019marabou], augmenting the training set whenever certificate constraints are violated. Such verification-oriented pipelines strengthen assurance but add computational overhead and face scalability challenges when certificates, neural policies, or state dimensions grow.
Lyapunov–barrier reward shaping
Finally, a smaller body of work explores shaping rewards using both CLF and CBF terms. Mizuta et al. [mizuta2024cobl] incorporate CLF/CBF rewards to guide diffusion-model denoising, illustrating a broader design space for reward shaping even though it is not purely within standard safe RL control benchmarks. In general, Lyapunov–barrier shaping remains comparatively under-explored in continuous-control safe RL, especially with hard safety objectives.
6.2 Key Issues and Open Problems
Feasibility and complexity of joint CLF–CBF constraints
Simultaneously enforcing Lyapunov decrease and barrier invariance can create infeasible QPs or induce conservative solutions (especially near boundaries), motivating adaptive relaxation strategies, feasibility certificates, and principled slack allocation [li2019temporal, meng2023reinforcement].
Learning certificates from data and guaranteeing generalization
Model-free CLBF/Lyapunov-like CBF critics rely on sampled transitions to learn certificates [du2023reinforcement, cohen2023safe]. Establishing conditions under which learned certificates remain valid under approximation error and distribution shift is still an open problem.
Closed-loop interactions between optimization filters and RL updates
When per-step QPs are used, the executed action differs from $u_{RL}$, which can bias gradients and off-policy learning, and complicate credit assignment. Differentiable QP layers and stability-aware policy gradient theory for filtered policies remain important directions.
Scalable verification and synthesis of neural certificates
Verification-driven pipelines (e.g., CEGIS) can significantly improve assurance [mandal2024formally], but scaling verification and counterexample generation to high-dimensional continuous control remains challenging.
Unified design principles for CLBF selection
Across works, CLF/CBF choices are often problem-specific. Developing systematic CLBF parameterizations (e.g., structured neural certificates, compositional certificates, reach-while-avoid certificates) and practical guidelines for selecting Lyapunov and barrier candidates is a key research opportunity.
Table 5 summarizes the literature reviewed for cumulative use of Lyapunov and barrier functions in RL. Overall, combined Lyapunov–barrier methods offer a compelling route to jointly guarantee stability and constraint satisfaction, but their broad deployment hinges on resolving feasibility/computation barriers for constrained optimization filters and establishing robust learning and verification principles for data-driven certificates.